OpenStudy (anonymous):

find f^-1(x) given f(x) =3sqrt(x-8)

OpenStudy (anonymous):

solve your f(x) function for x. then replace the f(x) with x and the x with f(x). If you need help just ask.

OpenStudy (whpalmer4):

\[y=3\sqrt{(x-8)}\]swap x and y, then solve for y \[x=3\sqrt{(y-8)}\] y=?

OpenStudy (anonymous):

sorry my internet was out for a couple minutes. ok, yes you can swap x and y at the beginning or the end of the process. x = 3*sqrt(y-8) x/3 = sqrt(y-8) (x/3)^2 = (y-8) (x/3)^2 + 8 = y hope that helps

OpenStudy (anonymous):

no, thanks I am talking about inverse not multiplying exponets, I have figured it out, thanks

OpenStudy (anonymous):

oh for anyone who make look up this question later the its f(-1)= 3x+8